Hi, I need help with this proportion.
4x/3x = 6x/4.5x
Solve for x.

Answer by checkley77(12844)

4x/3x = 6x/4.5x CROSS MULTIPLY
3*6X^2=4*4.5X^2
18X^2=18X^2
X^2=18X^2/18
X^2=X^2 INFINITE SOLUTIONS FOR X.